OpenStudy (anonymous):

If a = 3, b = -1, then 2a^2 - 2ab +b^2 = ?

OpenStudy (jlastino):

Just replace the given values into the equation

OpenStudy (anonymous):

Is the correct answer 25?

OpenStudy (jlastino):

yeah :D

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

\[(2 \times 3^2 ) -(2 \times 3 \times -1) + (-1)^2\] 18+6+1 = 25

OpenStudy (anonymous):

thanks